OpenStudy (cutiekawdfi):

HELP will fan and medal identify the conic section, lines of symmetry, and domain and range 1. x^2+y^2=4 2. x=2y^2+5

OpenStudy (briannaontimeforreal):

e^x/(e^(2x)+2e^x+1); y'(x)=y(x)+x+x y(x)+1,y(0)=1

OpenStudy (cutiekawdfi):

?

OpenStudy (88):

hi

OpenStudy (briannaontimeforreal):

thats the answer 42 briannaontimeforreal Medals 0 e^x/(e^(2x)+2e^x+1); y'(x)=y(x)+x+x y(x)+1,y(0)=1

OpenStudy (briannaontimeforreal):

hi 88

OpenStudy (cutiekawdfi):

which one

OpenStudy (briannaontimeforreal):

e^x/(e^(2x)+2e^x+1); y'(x)=y(x)+x+x y(x)+1,y(0)=1

OpenStudy (briannaontimeforreal):

that one

OpenStudy (cutiekawdfi):

which question